Determine the present value P you must invest to have the future value A at simple interest rate r after time t. A​ = ​$2000.00​, r​ = 9.0​%, t​ = 3 months
Answer #1

A = P(1 + rt)

Where A = amount = 2000

P = present value.

r = rate of interest = 0.09

t = time in year = 3/12

So,

2000 = P ( 1 + 0.09 * 3/12)

Solve for P

P = 1955.99